Output Total Cost 0 18 25 36 48 65 84 4 6 Refer to the table of output and total cost above. What is the total fixed cost when the output is 4? A) 2 c) $12 D) $40 E) $48 Help please
History Books QTotal Utility Q Total Utility Novels 0 0 0 180 1 140 300 260 -390 13 4 450 4360 5 480 5 380 390 495 Refer to the table above. The price of history books is $30 each and the price of novels is $20 each. Antonio has $200 of gift certificates to the local bookstore. He intends to purchase a combination of history books and novels. How many novels will he purchase? A) 1 B) 2 c) 3 D) 4
Given the information in the table, what is the economys Gross Domestic Product (GDP)? Amount (billions Item Consumption Income Investment Taxes Government Expenditures Transfer Payments Net Exports 12 2 3 3 -1 A) 28 billion B) 9 billion C) 16 billion D) 12 billion E) 27 billion
Answer #1

Q.1. Ans: B) $8

Explanation:

Fixed costs are those costs which are not changed with the change in the production process. So fixed cost is there even at zero level of output.

Q.2.Ans: D) 4

Explanation:

MUB / PB = MUN / PN

MU = Marginal utility

P = Price

Budget line (M) = ( $30 * 4 ) + ( $20 * 4 ) = $200

Q.3.Ans: C ) 16 billion

Explanation:

GDP = C + I + G + ( X - M )

= 12 + 2 + 3 - 1

= 17- 1

= 16

Where (X-M ) refers to net exports.